# Cryptomite: A versatile and user-friendly library of randomness extractors

### v1

@@ Abstract
We present Cryptomite, a Python library of randomness extractor implementations. The library offers a range of two-source, seeded and deterministic randomness extractors, together with parameter calculation modules, making it easy to use and suitable for a variety of applications. We also present theoretical results, including new extractor constructions and improvements to existing extractor parameters. The extractor implementations are efficient in practice and tolerate input sizes of up to $2^{40}>10^{12}$ bits. They are also numerically precise (implementing convolutions using the Number Theoretic Transform to avoid floating point arithmetic), making them well suited to cryptography. The algorithms and parameter calculation are described in detail, including illustrative code examples and performance benchmarking.
@@ 1 Introduction
Informally, a randomness extractor is a function that outputs (near-)perfect randomness (in the sense that it is almost uniformly distributed) when processing an input that is ‘somewhat’ random (in the sense that it may be far from uniformly distributed). Randomness extractors play an essential role in many cryptographic applications, for example as exposure-resilient functions, to perform privacy amplification in quantum key distribution or to distil cryptographic randomness from the raw output of an entropy source. Beyond cryptography, randomness extractors are a useful tool for many tasks, including the de-randomisation of algorithms or constructing list-decodable error-correcting codes.
Despite their usefulness, a major difficulty is to select the appropriate randomness extractor and its parameters for the task at hand – and more generally, to design, optimise and implement the algorithm without significant expertise and time investment. To address this challenge, we have developed Cryptomite, a software library that offers state-of-the-art randomness extractors that are easy to use (see Section 4: ‘Cryptomite: Examples with Code’), efficient (see Section 2.3: ‘Performance’) and numerically precise (we use the number theoretic transform (NTT) [1] whenever implementing a suitable convolution to avoid floating point arithmetic).
The library contains extractors from existing work and our own constructions, along with several improvements to parameters and security proofs. With the library, we provide explanations, theorems and parameter derivation modules to assist the user and ensure that they make a good extractor choice for their application. Cryptomite is made available at https://github.com/CQCL/cryptomite under a public licence https://github.com/CQCL/cryptomite/blob/main/LICENSE for non-commercial use, and installable through pip, with command pip install cryptomite. In [2], we additionally show how to use the extractors in Cryptomite to post-process the output of several (quantum) random number generators and study their effect using intense statistical testing. We have also developed extractors that have: a) been further optimised for performance, b) been implemented on Field Programmable Gate Arrays (FPGA) and c) been implemented in constant time to avoid possible side-channel attacks. The details can be found in Section 2.4.
@@ 1.1 Randomness Extraction at a Glance
Randomness extractors transform strings of weakly random numbers (which we call the source or input), in the sense that their distribution has some min-entropy (the most conservative measure of a random variable’s unpredictability, see Definition 1), into (near-)perfect randomness, in the sense that the output’s distribution is (almost) indistinguishable from uniform (see Definitions 3 and 4), as quantified by an extractor error $\epsilon$. Some extractors require an additional independent string of (near-)perfect random numbers for this transformation, which we call the seed. Additionally, we also call the seed weak if it is not (near-)perfect but only has min-entropy.
In the manuscript, we use the notation $n_{1},n_{2}$ to denote the lengths (in bits) of the weakly random extractor input and the (weak) seed respectively. We use $k_{1},k_{2}$ (respectively) to denote their min-entropy, $m$ the extractor’s output length, $\epsilon$ the extractor error and $O(.)$ the asymptotic quantities related to the extractor algorithm, for example its computation time. The set-up and variables are illustrated in Figure 1.
Figure 1: An extractor processes a weak input of $n_{1}$ bits with min-entropy $k_{1}$ in order to produce an output of $m$ bits that is $\epsilon$-perfectly random. Seeded and two-source extractors additionally require a second bit string called the (weak) seed, of $n_{2}$ bits and min-entropy $k_{2}$ (with $k_{2}=n_{2}$ for seeded extractors).
@@
  • Multi-Source Extractors, sometimes called blenders, are a generalisation of seeded extractors. They require one or multiple additional independent weak sources of randomness instead of a single (near-)perfect and independent one. In this work, the only multi-source extractors we consider are two-source, i.e. where the user has one addition weak input string only, which we call the weak seed.
@@
  • Implementability and efficiency: Some randomness extractors only have non-constructive proofs and are therefore theoretical objects that can only be used to derive fundamental results. Moreover, although a randomness extractor may have an explicit construction, this does not always mean that the algorithm can be implemented with a computation time suitable for the application at hand. For example, taking directly the building blocks in [4] would lead to a computation time of $O(n_{1}^{4})$ (for $n_{1}\propto n_{2}$), which significantly limits the maximum input lengths that the implementation can handle. For most applications, a computation time of $O(n_{1}^{2})$ is a requirement, whilst many even require quasi-linear computation time $O(n_{1}\log(n_{1}))$.
@@
  • Additional resources: Seeded and two-source extractors require a seed, which is either near-perfect (seeded extractors) or weak (two-source extractor). In addition to the weakly random input to extract from, this oblige the user to have access to, or be able to generate, an independent (weak) seed, which can be difficult to justify (and often not even discussed). Because of this, it is important to minimise the resources required by the extractor, such as for example two-source extractors with low entropy requirements on the weak seed [5, 6] or extractors that suffice with a (near-)perfect seed that is only short [7, 8].
@@
  • Security options: Certain randomness extractor constructions have been shown to fail against a quantum adversary who can store (side-)information about weak sources in quantum states [9]. Because of this, it is important to have the option to choose extractors which are quantum-proof, for example in protocols such as quantum key distribution. Similarly, in certain protocols the adversary is able to perform actions that may correlate the input and the (weak) seed, therefore breaking the independence condition. This is the case for example in randomness amplification or could happen in quantum key distribution. Because of this, techniques exist that allow to weaken the independence condition [10, 11, 12].
@@
  • Entropy loss: Entropy loss is the difference between the min-entropy of the extractor input and the length of the extractor output. Some randomness extractors are able to extract with entropy loss that is only logarithmic in the inverse extractor error, which is the fundamental minimum [13], against both quantum and classical adversaries.
@@
In Cryptomite, we implemented several randomness extractors that have a variety of the important features presented above. The extractors are efficient in practice, implemented with $O(n\log n)$ or $O(n^{2})$ time complexity, and have minimal or near-minimal entropy loss. They are able to extract from a wide array of different initial resources, allowing a user to pick the one most suited to their protocol. They also come with a wide range of options; such as whether to make them quantum-proof or secure when the extractor input and the (weak) seed are not fully independent. All extractors that we implement are strong, i.e. guarantee that the extractor output is statistically independent from one of its inputs, which for example allows for an increased output length by composing extractors together (see Section 3.2.5 and Definition 8 for the formal statement). They are information-theoretically secure, i.e. do not rely on computational assumption made on the adversary, and universally composable [14], allowing the extractor to be used securely as a part in a wider cryptographic protocol.
@@ 1.2 Our Contributions
    We provide an efficient code implementation of a two-source $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor, based on [15] and our previous work [16]. It is implemented in quasi-linear computation time $O(n\log n)$.
@@
    We present a new seeded extractor construction, which we call $\mathsf{Circulant}$, that can be understood as an extension of the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor. Beyond its simplicity, it is directly quantum-proof with the same error (and equivalently, the same output length) [^1] The new construction is now a two-universal family of hash functions, allowing to apply well-known security proof techniques, as such, the results in [17] prove it is quantum-proof with the same output length in the seeded case., making it the best choice against quantum adversaries. It achieves the same entropy loss, quasi-linear efficiency and error as Toeplitz-based ones [18], whilst requiring a seed that is the same length as the weak input only – making it an excellent choice for tasks such as quantum key distribution.
@@
    We implement the $\mathsf{Toeplitz}$ extractor [18] efficiently (in $O(n\log(n))$) and a $\mathsf{Trevisan}$ extractor, based on [7] and [19], which tolerates the shortest seed lengths for sufficiently large input lengths [^2] This is because the asymptotic claim that the seed length is $O(\log(n))$ does not mean that the seed length is short in practice. For example, when using the constructions of [19] that give near minimal entropy loss, the seed size is in reality larger than the input for input lengths below $10^{6}$ bits when outputting over $10^{4}$ bits with error less than $2^{-32}$. but is less efficient ($O(n^{2}\log(\log(n)))$ complexity, making it impractical in many cases). We also provide an implementation of the deterministic $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor in $O(n)$.
@@
    We implement the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors using the number theoretic transform (NTT) which gives a throughput of approximately $1$Mbit/s for input sizes up to $5\times 10^{8}$ on a standard personal laptop (see Section 2.3). Using the NTT instead of the fast Fourier transform (FFT) avoids numerical imprecision (due to floating point arithmetic) which may be unsuitable for some applications, e.g. in cryptography, especially with large input lengths. The software is able to process input lengths of up to $2^{40}>10^{12}$ bits, which should be sufficient even for device independent protocols. When using input lengths below $2^{29}\approx 5\cdot 10^{8}$, the code uses a further optimised (smaller finite field for the) NTT to increase the throughput.
@@
    We collate (and sometimes improve) results from existing works into a single place, providing techniques to get the most (near-)perfect randomness in different adversarial models and experimental settings. For example, all our constructions can, as an option, extract under a weaker independence requirement on the weak seed (in the Markov model [11]) and we provide the associated parameters. Moreover, in the software we include ḟrom_params utility functions that calculates the output and input lengths in a number of settings.
To the best of our knowledge, related works have some, but not all, of the above features. In particular, $\mathsf{Circulant}$ is new to this work and no alternative work uses the NTT for extractor implementation. In terms of other works that implement randomness extractors, in [19], the authors provide a publicly available implementation of the $\mathsf{Trevisan}$ extractor only, whilst [20] gives a library of several randomness extractors (including $\mathsf{Von\hskip 2.84544ptNeumann}$, $\mathsf{Toeplitz}$ and $\mathsf{Trevisan}$), but does not implement the $\mathsf{Toeplitz}$ extractor with quasi-linear computation time, i.e. does not use the FFT or NTT. There are several publicly available implementations of the $\mathsf{Toeplitz}$ extractor alone, including [21, 22], none of which use the NTT to achieve both quasi-linear computation time and numerical precision. In [23] the authors discuss and implement the modified $\mathsf{Toeplitz}$ extractor, using the FFT and based on [8], which is a seeded extractor with similar parameters to $\mathsf{Circulant}$. We give a more detailed comparison between the constructions in Section 3.5. Other works collate existing results on extractors with the focus on aiding a user to implement and select them for their protocol, for example [24], however they do not provide the same level of flexibility or options as this work.
@@ 2.1 Extractors of Cryptomite
The Cryptomite library contains
  • The (new) $\mathsf{Circulant}$ seeded extractor. It requires a prime seed length of $n_{2}=n_{1}+1$ and outputs $m=\lfloor k_{1}-2\log_{2}(\frac{1}{\epsilon})\rfloor$ bits against both classical and quantum adversaries (called classical- and quantum-proof respectively). We offer different two-source extensions of the extractor, one of which has output length $m=\lfloor k_{1}+k_{2}-n_{2}-2\log_{2}(\frac{1}{\epsilon})\rfloor$ against an adversary with quantum side-information that is product on the two sources (see Section 3.2.1), and the other outputs roughly a fifth of that in the stronger Markov model of side-information [11], which can be informally understood as a model in which the inputs are independent only when conditioned on the adversary’s side-information (see Section 3.2.2).
  • The $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ two-source extractor [15], with implementation based on [16]. It requires equal length inputs $n_{1}=n_{2}=n$, where $n$ is a prime with primitive root 2 and outputs $m=\lfloor k_{1}+k_{2}-n-2\log_{2}(\frac{1}{\epsilon})\rfloor$. Its quantum-proof version in the Markov model outputs $m=\lfloor\frac{1}{5}(k_{1}+k_{2}-n+8\log_{2}\epsilon+9-4\log_{2}(3))\rfloor$.
  • The $\mathsf{Toeplitz}$ strong seeded extractors based on [18]. It requires a seed length of $n_{2}=n_{1}+m-1$ and outputs $m=\lfloor k_{1}-2\log_{2}(\frac{1}{\epsilon})\rfloor$, when classical-proof or quantum-proof with the same output length. We also offer two-source extensions of this extractor, whereby, the output becomes $m=\lfloor\frac{1}{2}(k_{1}+k_{2}-n_{1}+1-2\log_{2}(\frac{1}{\epsilon}))\rfloor$ against product quantum side-information and a third of that in the Markov model.
  • The $\mathsf{Trevisan}$ seeded extractor [7], with implementation based on [19]. It requires a seed length of $n_{2}=O(\log(n_{1}))$ (see the exact statement of the seed length in Section 3.6) and outputs $m=\lfloor k_{1}-6-4\log_{2}(\frac{m}{\epsilon})\rfloor$ when classical-proof or quantum-proof with the same output length. As with $\mathsf{Circulant}$ and $\mathsf{Toeplitz}$, we offer two-source extensions, with the details deferred to Section 3.6.
@@
Our implementations of the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors all have $O(n_{1}\log(n_{1}))$ computation time and use the NTT. For these extractors, our code can tolerate input lengths $n_{1}\leq 2^{40}\approx 10^{12}$ bits – with an improved throughput version for any input lengths under $2^{29}$ (for more details, see Section 2.3). Our implementation of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor has $O(n_{1})$ computational complexity and $\mathsf{Trevisan}$ has $O\left(n_{1}^{2}\log\log\left(\frac{n_{1}^{2}}{\epsilon}\right)\right)$. Again, all two-source and seeded extractors mentioned above are strong (see Definition 8) with the same error and output length.

### v3

@@ Abstract
We present Cryptomite, a Python library of randomness extractor implementations. The library offers a range of two-source, seeded and deterministic randomness extractors, together with parameter calculation modules, making it easy to use and suitable for a variety of applications. We also present theoretical results, including new extractor constructions and improvements to existing extractor parameters. The extractor implementations are efficient in practice and tolerate input sizes of up to $2^{40}>10^{12}$ bits. Contrary to alternatives using the fast Fourier transform, we implement convolutions efficiently using the number-theoretic transform to avoid rounding errors, making them well suited to cryptography. The algorithms and parameter calculation are described in detail, including illustrative code examples and performance benchmarking.
@@ 1 Introduction
Informally, a randomness extractor is a function that outputs (near-)perfect randomness (in the sense that it is almost uniformly distributed) when processing an input that is ‘somewhat’ random (in the sense that it may be far from uniformly distributed). Randomness extractors play an essential role in many cryptographic applications, for example as exposure-resilient functions, to perform privacy amplification in quantum key distribution or to distil cryptographic randomness from the raw output of an entropy source. Beyond cryptography, randomness extractors are a useful tool for many tasks, including the de-randomisation of algorithms or constructing list-decodable error-correcting codes. For an introduction to randomness extraction, see [1].
Despite their usefulness, a major difficulty is to select the appropriate randomness extractor and its parameters for the task at hand – and more generally, to design, optimise and implement the algorithm without significant expertise and time investment. To address this challenge, we have developed Cryptomite, a software library that offers state-of-the-art randomness extractors that are easy to use (see Section 5: ‘Cryptomite: Examples with Code’) and efficient (see Section 2.3: ‘Performance’). Our implementations are also numerically precise, in the sense that the extractor algorithms do not use any floating point arithmetic that can lead to rounding errors. This is accomplished using the number-theoretic transform (NTT) [2], instead of the fast Fourier transform (FFT) [3], whenever implementing convolutions efficiently. [^1] The NTT can be seen as a generalisation of the FFT to finite fields, allowing to perform fast convolutions on integer sequences. The FFT directly performs fast convolutions using floating point numbers and thus precision limited by the amount of memory dedicated to it. This is important since uncontrolled errors can lead to implementation vulnerabilities or failures, making our implementations more reliable than alternatives.
The library contains extractors from existing work and our own constructions, along with several improvements to parameters and security proofs. With the library, we provide explanations, theorems and parameter derivation modules to assist the user and ensure that they make a good extractor choice for their application. Cryptomite is made available at https://github.com/CQCL/cryptomite under a public licence https://github.com/CQCL/cryptomite/blob/main/LICENSE for non-commercial use, and installable through pip, with command pip install cryptomite. In [4], we additionally show how to use the extractors in Cryptomite to post-process the output of several (quantum) random number generators and study their effect using intense statistical testing.
@@ 1.1 Randomness extraction at a glance
Randomness extractors transform strings of weakly random numbers (which we call the source or input), in the sense that their distribution has some min-entropy (the most conservative measure of a random variable’s unpredictability, see Definition 1), into (near-)perfect randomness, in the sense that the output’s distribution is (near-)indistinguishable from uniform (see Definitions 3 and 4), as quantified by an extractor error $\epsilon$. Some extractors require an additional independent string of (near-)perfect random numbers for this transformation, which we call the seed [^2] Remark that, for useful seeded extractors, the seed is either short or the extractor is strong (which means that its output is (almost) independent of the seed and therefore the seed is not consumed in the process, see Definition 8 for the formal statement).. Additionally, we call the seed a weak seed if it is not (near-)perfect and has only a min-entropy guarantee.
In the manuscript, we use the notation $n_{1},n_{2}$ to denote the lengths (in bits) of the weakly random extractor input and the (weak) seed respectively. We use $k_{1},k_{2}$ (respectively) to denote their min-entropy, $m$ the extractor’s output length, $\epsilon$ the extractor error and $O(\cdot)$ the asymptotic quantities related to the extractor algorithm, for example its computation time. The set-up and variables are illustrated in Figure 1.
Figure 1: An extractor processes a weak input of $n_{1}$ bits with min-entropy $k_{1}$ in order to produce an output of $m$ bits that is $\epsilon$-perfectly random. Seeded and two-source extractors additionally require a second random bit string called the (weak) seed, of $n_{2}$ bits and min-entropy $k_{2}$ (with $k_{2}=n_{2}$ for seeded extractors).
@@
  • Multi-source extractors, sometimes called blenders, are a generalisation of seeded extractors. They require at least one additional independent weak source of randomness, rather than a single (near-)perfect and independent one. In this work, the only multi-source extractors we consider are two-source, i.e. where the user has one addition weak input string only, which we call the weak seed.
@@
  • Implementability and efficiency: Some randomness extractors only have non-constructive proofs and are therefore theoretical objects that can only be used to derive fundamental results. Moreover, although a randomness extractor may have an explicit construction, this does not always mean that the algorithm can be implemented with a computation time suitable for the application at hand. For example, taking directly the building blocks in [6] would lead to a computation time of $O(n_{1}^{4})$ (for $n_{1}\propto n_{2}$), which significantly limits the maximum input lengths that the implementation can handle. For most applications, a computation time of $O(n_{1}^{2})$ is a requirement, whilst many even require quasi-linear computation time $O(n_{1}\log(n_{1}))$ – such as quantum key distribution. For an excellent discussion and concrete example of the need for quasi-linear computation time, see Appendix E, Section C of [7] and our examples with code in Section 5.
@@
  • Additional resources: Seeded and two-source extractors require a seed, which is either near-perfect (seeded extractors) or weak (two-source extractor). In addition to the weakly random input to extract from, this obliges the user to have access to, or be able to generate, an independent (weak) seed, which can be difficult to justify (and often not even discussed). Because of this, it is important to minimise the resources required by the extractor, for example, by extracting with a (near-)perfect seed that is only short [8, 7] or minimising the entropy requirement on the weak seed [9, 10].
@@
  • Security options: Certain randomness extractor constructions have been shown to fail against a quantum adversary who can store information about extractor inputs in quantum states [11]. Because of this, it is important to have the option to choose extractors that are quantum-proof, for example, in protocols such as quantum key distribution. Similarly, in certain protocols the adversary is able to perform actions that may correlate the input and the (weak) seed, therefore breaking the independence condition. This is the case for example in randomness amplification or could happen in quantum key distribution. Consequently, techniques have been developed that allow to weaken the independence condition [12, 13, 14].
@@
  • Entropy loss: Entropy loss is the difference between the min-entropy of the extractor input and the length of the extractor output. Some randomness extractors are able to extract with entropy loss that is only logarithmic in the inverse extractor error, which is the fundamental minimum [15] against both quantum and classical adversaries.
@@
In Cryptomite, we implemented several randomness extractors that have a variety of the important features presented above. Our extractors are efficient, with $O(n_{1}\log n_{1})$ or $O(n_{1}^{2}\log^{2}n_{1})$ computation time, and have minimal or near-minimal entropy loss. They can extract from various initial resources and come with many options; such as whether to make them quantum-proof or secure when the input and (weak) seed are not fully independent. Additionally, all of our extractors are strong (meaning the output is statistically independent of one input), information-theoretically secure (secure against computationally unbounded adversaries) and universally composable [16] (enabling secure integration into broader cryptographic protocols).
@@ 1.2 Our contributions
    We provide an efficient code implementation of a two-source $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor, based on [17] and our previous work [18]. It is implemented in quasi-linear computation time $O(n_{1}\log n_{1})$. To our knowledge, there exist no alternative implementation.
@@
    We present a new strong seeded extractor construction, which we call $\mathsf{Circulant}$, due to the input being used to construct a circulant matrix. Beyond its simplicity, it is directly quantum-proof with the same error (and equivalently, the same output length) [^3] The new construction is now a two-universal family of hash functions, allowing to apply well-known security proof techniques, as such, the results in [19] prove it is quantum-proof with the same output length in the seeded case., making it the best choice against quantum adversaries. It achieves the same entropy loss, quasi-linear computation time and error as Toeplitz-based ones [20], whilst requiring a seed that is only a single bit longer than the weak input length – making it an excellent choice for tasks such as quantum key distribution. This new extractor construction also has no alternative implementation.
@@
    We implement the $\mathsf{Toeplitz}$ extractor [20] efficiently (in $O(n_{1}\log(n_{1}))$) and a $\mathsf{Trevisan}$ extractor, based on [8] and [21], which tolerates the shortest seed lengths for sufficiently large input lengths but is less efficient ($O(mn_{1}\log^{2}n_{1})$ computation time, where $m$ is the output size of the extractor – making it impractical in many cases) [^4] We note that the seed length is $O(\log(n_{1}))$, but this does not always imply a short seed in practice. For instance, in our implementation with near-minimal entropy loss, the seed length exceeds the input length when $n_{1}<10^{6}$ and $m\approx n_{1}/2$. Furthermore, it is often useful to assume that the output length $m$ is proportional to the input length $n_{1}$, i.e., $m\propto n_{1}$. In this case, the computation time for the Trevisan extractor becomes $O(n_{1}^{2}\log^{2}n_{1})$, making it inefficient. However, when $m$ is small, Trevisan’s extractor can have a reasonable computation time, as seen in [22], where $m=512$. We also provide an implementation of the deterministic $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor in $O(n_{1})$.
@@
    Contrary to all alternative implementations, we implement the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors using the number-theoretic transform (NTT) implemented in C++, which gives a throughput of approximately 11Mbit/s for input sizes up to $5\times 10^{8}$ on a standard personal laptop (see Section 2.3). Using the NTT instead of the fast Fourier transform (FFT) avoids numerical imprecision (due to floating point arithmetic) which may be unsuitable for some applications, e.g. in cryptography, especially with large input lengths. Contrary to alternatives, our software is able to process input lengths of up to $2^{40}>10^{12}$ bits, which should be sufficient even for device-independent protocols [23, 24, 25, 26, 18, 27, 28]. When using input lengths below $2^{29}\approx 5\cdot 10^{8}$, the code uses a further optimised (smaller finite field for the) NTT to increase the throughput.
@@
    We collate (and sometimes improve, see for example Corollary 1) results from existing works into a single place, providing techniques to get the most (near-)perfect randomness in different adversarial models and experimental settings. For example, all our constructions can, as an option, extract under a weaker independence requirement on the weak seed (in the Markov model [13]) and we provide the associated parameters. Moreover, in the software we include ḟrom_params utility functions which calculate the input and output lengths for a number of settings and a suggest_extractor function, which assists a user on deciding which extractor to use for their application, based on the flow chart in Figure 2.
To the best of our knowledge, existing works only achieve a small subset of the features mentioned above. There are no alternative implementations of the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor or the $\mathsf{Circulant}$ extractor, which are new to this work. Reference [29] is the only work we found that offers a library containing several software implementations of randomness extractors, including the $\mathsf{Von\hskip 2.84544ptNeumann}$ and $\mathsf{Toeplitz}$ extractors, as well as one termed ‘universal hashing’ and providing a link to the implementation of the $\mathsf{Trevisan}$ extractor [30]. However, it does not implement the $\mathsf{Toeplitz}$ extractor efficiently by exploiting the convolution theorem to obtain quasi-linear computation time, i.e., it does not use the FFT or NTT.
Other software implementations of individual extractors exist. There are several publicly available implementations of the $\mathsf{Toeplitz}$ extractor, for example, in [31, 32] (in Python) and [33] (in C++ and Verilog for FPGA), but none utilise the convolution theorem for quasi-linear computation time – making them inefficient at larger input sizes. Implementations of the $\mathsf{Trevisan}$ extractor can be found in [34], related to [21] and [30], both written in C++. The implementation of [34] is particularly flexible, providing various instantiations of the extractor that can be useful to experts. In [35], the authors discuss and implement a modified $\mathsf{Toeplitz}$ extractor, based on [7], which is an extractor with parameters similar to $\mathsf{Circulant}$. This implementation uses the FFT, enabling input lengths up to $10^{7}$. However, by not using the NTT as we do, potential rounding errors may occur. A more detailed comparison is provided in Section 4.3.
@@ 2.1 Extractors of Cryptomite
The Cryptomite library contains:
  • The (new) $\mathsf{Circulant}$ strong seeded extractor. It requires a prime with primitive root 2 seed length, $n_{2}=n_{1}+1$, and outputs $m=\lfloor k_{1}-2\log_{2}(\frac{1}{\epsilon})\rfloor$ bits against both classical and quantum adversaries (called classical- and quantum-proof respectively). We present a two-source extension of this extractor, secure against quantum side information that is either product across the two sources (see Section 3.2.1) or follows the Markov model of side information [13] (see Section 3.2.2), with output length $m=\lfloor k_{1}+(k_{2}-n_{2})-2\log_{2}(\frac{1}{\epsilon})\rfloor$.
  • The $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ strong two-source extractor [17], with implementation based on [18]. It requires equal length inputs $n_{1}=n_{2}=n$, where $n$ is a prime with primitive root 2 and outputs $m=\lfloor k_{1}+k_{2}-n-2\log_{2}(\frac{1}{\epsilon})\rfloor$. Its quantum-proof version in the Markov model outputs $m=\lfloor\frac{1}{5}(k_{1}+k_{2}-n+8\log_{2}\epsilon+9-4\log_{2}(3))\rfloor$.
  • The $\mathsf{Toeplitz}$ strong seeded extractors based on [20]. It requires a seed length of $n_{2}=n_{1}+m-1$ and outputs $m=\lfloor k_{1}-2\log_{2}(\frac{1}{\epsilon})\rfloor$, when classical-proof or quantum-proof with the same output length. We also present a two-source extension of this extractor, secure against quantum side information that is either product across the two sources or follows the Markov model of side information, with output length $m=\lfloor k_{1}+(k_{2}-n_{2})-2\log_{2}(\frac{1}{\epsilon})\rfloor=\lfloor\frac{1}{2}(k_{1}+k_{2}-n_{1}+1-2\log_{2}(\frac{1}{\epsilon}))\rfloor$.
  • The $\mathsf{Trevisan}$ strong seeded extractor [8], with implementation based on [21]. It requires a seed length of $n_{2}=O(\log(n_{1}))$ (see the exact statement of the seed length in Section 4.4) and outputs $m=\lfloor k_{1}-6-4\log_{2}(\frac{m}{\epsilon})\rfloor$ when classical-proof or quantum-proof with the same output length. As with $\mathsf{Circulant}$ and $\mathsf{Toeplitz}$, we offer two-source extensions, with the details deferred to Section 4.4.
@@
Our implementations of the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors all have $O(n_{1}\log(n_{1}))$ computation time and use the NTT. For these extractors, our code can tolerate input lengths $n_{1}\leq 2^{40}\approx 10^{12}$ bits – with an improved throughput version for any input lengths under $2^{29}$ (for more details, see Section 2.3). Our implementation of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor has a computation time of $O(n_{1})$, while $\mathsf{Trevisan}$ has a computation time of $O\left(n_{1}m\log^{2}\left(\frac{n_{1}m}{\epsilon}\right)\right)$.